A certain mass of hydrogen gas collected over water at 6degrees Celsius and 765mmHg pressure has a volume of 35cm3. Calculate the volume when it is dry at STP
[SVP of water at 6degrees Celsius = 7mmHg]
1
Expert's answer
2020-04-24T13:36:05-0400

"P_{total} = P_{H_2}+P_{H_2O}"

"765=P_{H_2}+ 7"

"P_{H_2} = 758 mmHg"

Use Combined Gas Law

"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"


"P_1 = 758 mmHg"

"V_1= 0.035 L"

"T_1 (K) = 273+6 = 279 K"


"P_2 = 760 mmHg" (pressure at STP)

"V_2=?"

"T_2=273 K" (temperature at STP)


"\frac{758\times0.035}{279}=\frac{760\times V_2}{273}"


"V_2 = 0.034 L"